European Science Foundation European Research Conferences

The Programme of European Research Conferences is run by the European Science Foundation, in association with several learned societies, and with funding from the Human Capital and Mobility Programme of the European Union. Each Conference consists of a series of meetings, held typically every other year. There are neither written contributions nor proceedings.

The Commission of the European Community subsidises EURESCO. Young mathematicians under the age of 35 who are EU citizens (or residents in an EU country for more than a year) are encouraged to apply to the EURESCO programme. The applications should be sent directly to the EURESCO office:

Dr. J. Hendekovic, EURESCO; European Science Foundation; 1, quai Lezay–Marnésia; F-67080 Strasbourg; France

I. MATHEMATICAL METHODS IN INDUSTRIAL PROBLEMS

1994 Multi-scale Analysis in Image Processing. Chair: Pierre-Louis Lions (Paris), Vice-Chair: Alfio Quarteroni (Milano) 15–20 October 1994. Lunteren, The Netherlands

1996: Mathematical and Numerical Models for the Simulation of Turbulent and Reactive Flows.

Chair: Alfio Quarteroni (Milano)

16–21 February 1996

Mont St. Odile (near Strasbourg),France

II. ALGEBRA AND DISCRETE

MATHEMATICS

1996:

Group Theory: Finite to Infinite.

Chair: Laszlo Babai (Chicago);

Vice-Chair: Anders Björner (Stockholm)

13–18 July 1996

Castelvecchio, Italy

III. MATHEMATICAL ANALYSIS

Local singularities of solutions to nonlinear and singular PDE’s

Chair: Bogdan Ziemian (Warsaw);

Vice-Chair: Werner Müller, MPI Bonn

September 19–24, 1995.

San Feliu de Guixols, Spain

IV. NUMBER THEORY AND

ARITHMETICAL GEOMETRY

The first conference will be organised by Gerhard Frey (Essen), and will take place in the fall of 1997.